# Hierarchical Multi-Agent MCTS for Safety-Critical Coordination in Mixed-Autonomy Roundabouts

### v3

@@ I INTRODUCTION
Navigation at roundabouts presents unique challenges for Connected and Autonomous Vehicle (CAVs) [1], particularly in mixed traffic environments where both CAVs and human-driven vehicles (HDVs) must safely coordinate their movements through complex circular geometries and multiple merging points [2, 3]. The complexity arises from the need to handle multiple types of critical interactions simultaneously while ensuring both safety and efficiency in roundabouts, a highly dynamic environment characterized by continuous merging, lane-changing, and exit decision behaviors [4]. This challenge becomes more pronounced as the interaction patterns among vehicles become more intricate due to the unique geometric constraints and yielding rules of roundabout environments. Therefore, a comprehensive understanding of both deterministic CAV behaviors and uncertain HDV behaviors is required [5, 6].
Traditional approaches to roundabout management often rely on rule-based decision-making methods, which attempt to generate conflict-free navigation sequences through preset yielding regulations and lane assignment strategies [7, 8]. These methods struggle to capture the complex and diverse decision-making behaviors of human drivers, where yielding decisions involve significant behavioral variability. The conventional priority-based strategies [9] ensure safety by enforcing strict yielding rules. However, some priority-based methods significantly increase the requirement of onboard communication quality. More sophisticated rule-based approaches incorporating machine learning and static- and dynamic-constraint-based optimization have been proposed, but their effectiveness has only been verified in simple, single-lane roundabouts [10].
The application of machine learning to roundabout navigation has evolved significantly, yet critical gaps remain between theoretical advances and practical deployment [11, 12]. While deep reinforcement learning methods, including multi-agent deep deterministic policy gradient approaches, have demonstrated success in controlled simulations [13, 14, 15], their performance degrades substantially when faced with the geometric diversity and behavioral uncertainty characteristic of real-world roundabouts [16]. The core challenge lies not in the learning capability itself, but in the fundamental mismatch between the assumptions underlying these methods and the requirements of safety-critical traffic coordination [17, 18, 19]. Specifically, neural network-based approaches suffer from three critical limitations in roundabout applications [20]. First, the exponential growth of the state-action space with increasing vehicle density makes it computationally prohibitive to explore all safety-critical scenarios during training [11, 12]. Second, the lack of interpretability prevents real-time safety verification, a non-negotiable requirement for mixed-autonomy deployments [13, 14]. Third, the sim-to-real gap is particularly pronounced in roundabouts due to the complex interplay between lane-changing decisions and continuous control, leading to unexpected behaviors when policies trained in simulation encounter novel (out-of-training-scope) traffic patterns [15, 17, 18, 19].
@@
This paper introduces a safety-critical multi-agent MCTS framework for coordinating mixed traffic at dual-lane roundabouts. The framework addresses the unique challenges posed by roundabout geometry, including lane-specific uncertainty modeling, exit proximity effects, and the complex interactions between inner and outer lane vehicles.
The main contributions are summarized as follows:
  • We propose a safety-critical multi-agent MCTS framework specifically designed for roundabout environments that integrates deterministic and probabilistic vehicle behavior predictions, enabling cooperative decision making among CAVs and HDVs in dual-lane circular traffic flows.
  • We develop a hierarchical safety assessment mechanism that systematically handles three critical interaction types in roundabout scenarios:CAV-to-CAV (C2C) [40], CAV-to-HDV (C2H) [41] and CAV-to-Boundary (C2B) [42], by using dynamic safety thresholds and spatiotemporal risk metrics tailored to circular geometry constraints.
  • We adopt an adaptive human driving behavior prediction framework that combines the deterministic Intelligent Driver Model (IDM) with lane-specific probabilistic distributions and exit proximity effects to effectively capture the heightened uncertainties in roundabout HDV behaviors, ensuring robust robot decision making in mixed circular traffic environments.
The rest of this paper is organized as follows: Section II describes the problem and system structure, Section III illustrates the safety-critical decision making system formulation, Section IV introduces the multi-agent MCTS process, Section V presents simulation results, and Section VI concludes the paper.
II Overview of the Proposed Safety-Critical Decision-Making Framework
We focus on a challenging scenario involving unsignalized, dual-lane roundabouts, where CAVs must navigate in the presence of both other CAVs and HDVs, all without the aid of traffic signals. The task requires accounting for both predictable autonomous behaviors and the inherent uncertainty in human driving patterns.
@@ II Overview of the Proposed Safety-Critical Decision-Making Framework
An overview of our proposed framework is depicted in Fig. 1. The architecture is composed of four key modules. First, the multi-agent MDP model is established, including formal definitions of states, actions, and dynamic models with embedded safety constraints (Sec. III-A). Next, we introduce a hierarchical safety mechanism that considers layered safety constraints across vehicle types (CAVs and HDVs) and environmental boundaries, using dynamically updated safety thresholds and predictive risk assessment for HDVs (Sec. III-B).
Building on this foundation, we develop a safety-aware multi-agent MCTS algorithm. This component includes the design of safety-encoded tree node structures, an exploration-exploitation balancing strategy using Upper Confidence Bound (UCB), and a policy extraction process that leverages rollout-based evaluations and backpropagation (Sec. IV-A). Finally, we formulate a multi-objective reward function tailored to simultaneously consider A2A, A2H, and A2R safety, along with motion efficiency and feasibility under dynamic constraints. This leads to an integrated reward optimization scheme (Sec. IV-B).
Fig. 2 illustrates the interactive environment surrounding a CAV navigating a roundabout. The figure highlights key spatial regions that shape decision-making, including a central interaction zone where vehicles must account for potential conflicts and trajectory overlaps, and a broader approach area where speed and heading adjustments begin in anticipation of entering the roundabout. The inner circle represents the main circulation area of the roundabout, while the shaded zones near the exits denote areas of heightened interaction, where behaviors such as yielding, merging, and exiting are most critical. These regions collectively define the spatial context in which CAVs must coordinate with surrounding agents to ensure safe and efficient navigation.
III Safety-Critical Decision Making System Formulation
@@ III-A Multi-Agent MDP Formulation for Dual-Lane Roundabout
Figure 2: Illustration of the interaction scenario.
The roundabout navigation decisions must also satisfy constraints over $T$:
@@
TABLE I: Safety distance adjustment factors
@@ III-B Hierarchical Safety Assessment for Roundabout Navigation
2) C2C and C2H safety assessment: The adaptive safety distance accounts for multiple risk factors via a base value and adjustment terms, as shown in Table I:
@@
where $d_{\min}$ is the minimum base distance, $\kappa_{v}$ is the velocity scaling factor, $\Delta v_{ij}=v_{i}-v_{j}$ is the relative velocity, $\Delta\phi_{ij}$ is the relative heading angle, $l_{i},l_{j}\in\{0,1\}$ denote lane indices, $v_{\text{ref}}$ is the reference velocity for normalization, $\beta_{k}$ are weighting parameters, and $\mathbb{1}_{\Omega}(\cdot)$ is the indicator function for vehicles in zone $\Omega$. $\mathcal{K}=\{v,\phi,l,z\}$ indexes the adjustment factors in Table I. Fig. 3 visualizes our safety-critical risk assessment framework for CAVs and HDVs navigating roundabouts. Fig. 3(a) presents the foundational distance-based risk function, characterized by an exponential decay curve, which quantifies collision urgency and informs immediate avoidance actions in MCTS decisions. The clearly defined safe distance threshold enables effective and computationally efficient binary safety decisions. Fig. 3(b) provides a spatial risk heat map for a representative roundabout scenario, highlighting high-risk zones at lane-changing and exit-merging points.
@@
3) Lane-Specific HDV Uncertainty Modeling: For C2H interactions, we develop a novel lane-specific uncertainty model that accounts for the distinct behavioral patterns in inner and outer lanes. The uncertainty model is formulated as:
@@
where $\psi(\hat{s}_{i},\hat{s}_{h})=\mathbb{I}(d(\hat{s}_{i},\hat{s}_{h})<d{\text{safe}}(s_{i},s_{h}))$ is the collision indicator function. The overall safety level for C2H interactions is:
@@ IV-A Multi-Agent MCTS for Roundabout Navigation
Figure 5: Illustration of the safety-critical MCTS framework. Each node stores the visit count $N_{n}$, value estimate $Q_{n}$, and UCB score. Nodes identified as unsafe (shown in purple) are pruned during the safety validation stage. The green path indicates the backpropagation process.
@@
where $e_{\text{exp}}$ is the exploration constant that balances exploitation and exploration, and $N_{p_{n}}$ is the visit count of the parent node. Fig. 5 illustrates the structure of a MCTS guided by the UCB strategy. Leaf nodes beneath the child nodes represent terminal or intermediate states, with one explicitly marked as an “Unsafe Node” to signify risky or undesirable outcomes according constrains.
@@ IV-D Computational Complexity Analysis
The computational complexity of the proposed multi-agent MCTS-based planning algorithm arises from three main components: joint-action tree expansion, safety validation, and rollout simulation.
Tree expansion considers joint actions over $N$ CAVs, each with $|\mathcal{A}|$ possible actions, resulting in a branching factor of $|\mathcal{A}|^{N}$. At each expanded node, safety validation involves pairwise CAV interactions ($O(N^{2})$), CAV-to-HDV checks ($O(NM)$), and lane-level uncertainty evaluations, typically contributing $O(NM\cdot d^{2})$ complexity, where $d$ is the state dimension.
@@
(a) Illustration of MCTS simulation at a signal-free, dual-lane roundabout.
@@ V EXPERIMENTAL EVALUATION
(a) Illustration of MCTS simulation at a signal-free, dual-lane roundabout.
@@ V-A Case 1: Dual-lane Roundabout (ROP = 100%)
Methods Average Arrive Average Collision Average Simulation
Rate (%) Rate (%) Time (s)
Stackelberg $75.6\pm 4.3$ $16.8\pm 5.0$ $33.1\pm 7.0$
Nash $82.3\pm 2.5$ $11.4\pm 3.8$ $45.6\pm 9.5$
Baseline $84.1\pm 2.6$ $13.7\pm 5.1$ $\mathbf{21.5}\pm 4.0$
$d_{\max}=3$ $90.2\pm 3.2$ $3.5\pm 2.2$ $25.6\pm\mathbf{3.1}$
$d_{\max}=8$ $\mathbf{94.4\pm 2.0}$ $\mathbf{0.2\pm 0.3}$ $26.5\pm 3.8$
@@
Fig. 7(b) presents a quantitative assessment of safety performance based on Post-Encroachment Time (PET) distributions. The proposed method with a maximum tree depth of $d_{\text{max}}=8$ delivers the best safety outcomes, with no observed instances of PET falling below the critical threshold of $1.0$ second. In comparison, the baseline approach exhibits $8.8\%$ violations, while the Stackelberg and Nash methods perform considerably worse, with violation rates of $26.1\%$ and $13.2\%$, respectively. These results underscore the effectiveness of our framework, which leverages dynamic risk evaluation and adaptive safety constraints to proactively manage both current and anticipated vehicle interactions.
Table II highlights the superior performance of our proposed approach with $d_{\text{max}}=8$, which achieves the highest arrival rate of $94.4\pm 2.0\%$ and an almost perfect safety record with only $0.2\pm 0.3\%$ collision rate. Although its computation time ($26.5\pm 3.8$ s) is slightly longer than that of the baseline method ($21.5\pm 4.0$ s), the substantial safety gain justifies this modest increase. In contrast, conventional approaches such as Stackelberg and Nash still suffer from higher collision rates of $16.8\%$ and $11.4\%$, respectively, demonstrating their limitations in handling complex interaction scenarios.
@@
Methods Average Arrive Average Collision Average Simulation
Rate (%) Rate (%) Time (s)
Stackelberg $54.8\pm 5.1$ $33.0\pm 6.0$ $42.3\pm 7.2$
Nash $60.2\pm 4.0$ $28.5\pm 5.2$ $57.1\pm 6.4$
Baseline $68.4\pm 5.0$ $16.5\pm 7.9$ $\mathbf{25.1\pm 3.5}$
$d_{\max}=3$ $86.1\pm 4.5$ $4.8\pm 2.6$ $27.5\pm 5.2$
$d_{\max}=8$ $\mathbf{92.9\pm 3.2}$ $\mathbf{1.5\pm 0.5}$ $30.1\pm 5.0$

### v4

@@ I INTRODUCTION
Navigation at roundabouts presents unique challenges for Connected and Autonomous Vehicles (CAVs) [1], particularly in mixed traffic where both CAVs and human-driven vehicles (HDVs) must safely coordinate through complex circular geometries and multiple merging points [42, 12]. The difficulty lies in managing several critical interactions simultaneously while ensuring safety and efficiency in a highly dynamic setting with continuous merging, lane-changing, and exit decisions [30]. This challenge is further amplified by intricate interaction patterns induced by roundabout-specific geometric constraints and yielding rules. Therefore, a comprehensive understanding of both deterministic CAV behaviors and uncertain HDV behaviors is essential [3, 29].
Traditional approaches to roundabout management often rely on rule-based decision-making methods, which generate conflict-free navigation sequences through preset yielding regulations and lane assignment strategies [8, 32]. These methods struggle to capture the complex decision-making behaviors of human drivers, where yielding involves significant behavioral variability. Conventional priority-based strategies [2] ensure safety by enforcing strict yielding rules, but may substantially increase the requirement for onboard communication quality. More sophisticated rule-based approaches incorporating machine learning and static- and dynamic-constraint-based optimization have been proposed, yet their effectiveness has only been verified in simple single-lane roundabouts [14].
The application of machine learning to roundabout navigation has advanced significantly, yet critical gaps remain between theory and practical deployment [24, 27]. While deep reinforcement learning methods, including multi-agent deep deterministic policy gradient approaches, have shown success in controlled simulations [28, 23, 7], their performance degrades when faced with the geometric diversity and behavioral uncertainty of real-world roundabouts [34]. The core challenge lies not in learning capability itself, but in the mismatch between underlying assumptions and the requirements of safety-critical traffic coordination [11, 43, 20]. Specifically, neural network-based approaches exhibit three key limitations [35]. First, the exponential growth of the state-action space with vehicle density makes it computationally prohibitive to cover all safety-critical scenarios during training [24, 27]. Second, limited interpretability prevents real-time safety verification, a non-negotiable requirement in mixed autonomy [28, 7]. Third, the sim-to-real gap is severe in roundabouts due to the interplay between lane-changing decisions and continuous control, leading to unexpected behaviors when policies encounter novel traffic patterns.
@@
This paper introduces a safety-critical multi-agent MCTS framework for coordinating mixed traffic at dual-lane roundabouts. The framework addresses the unique challenges posed by roundabout geometry, including lane-specific uncertainty modeling, exit proximity effects, and complex interactions between inner and outer lane vehicles. Unlike prior MCTS-based planners that apply uniform safety checks or treat all traffic participants identically, our approach introduces a structurally differentiated design: safety assessment, uncertainty modeling, and reward allocation are tailored to the specific interaction type (CAV-to-CAV (C2C) [18], CAV-to-HDV (C2H) [13] and CAV-to-Boundary (C2B) [33]) and lane context, enabling these components to reinforce each other through hierarchical integration rather than operating as independent modules. The main contributions are summarized as follows:
  • We propose a safety-critical multi-agent MCTS framework for dual-lane roundabouts, unifying safety pruning, uncertainty modeling, and reward shaping into an interaction-aware tree search pipeline for mixed traffic.
  • We develop a structurally differentiated safety assessment mechanism that handles various interactions via lane-dependent risk metrics and conservative thresholds for human uncertainty, validated through ablation studies.
  • We introduce two key structural components for robust coordination: (i) an adaptive HDV prediction model capturing lane-specific uncertainty and exit proximity effects, and (ii) a Shapley value-based reward allocation scheme enabling contribution-aware credit assignment and reducing collision rates in cooperative roundabout navigation.
II Safety-Critical Decision-Making Framework
We consider unsignalized dual-lane roundabouts, where CAVs must navigate among other CAVs and HDVs without traffic signals. This task requires handling both predictable autonomous behaviors and uncertainty in human driving patterns.
@@ II Safety-Critical Decision-Making Framework
An overview of our proposed framework is depicted in Fig. 1. The architecture is composed of four key modules. First, the multi-agent MDP model is established, including formal definitions of states, actions, and dynamic models with embedded safety constraints (Sec. III-A). Next, we introduce a hierarchical safety mechanism that considers layered safety constraints across vehicle types and environmental boundaries, using dynamically updated safety thresholds and predictive risk assessment for HDVs (Sec. III-B).
Figure 2: Illustration of the interaction scenario.
Building on this foundation, we develop a safety-aware multi-agent MCTS algorithm with safety-encoded tree nodes, UCB-based exploration–exploitation balancing, and rollout-based policy extraction with backpropagation (Sec. IV-A). Finally, we formulate a multi-objective reward function that jointly accounts for A2A, A2H, and A2R safety, motion efficiency, and dynamic feasibility, yielding an integrated optimization scheme (Sec. IV-B).
Fig. 2 illustrates the interactive environment surrounding a CAV navigating a roundabout. The figure highlights key spatial regions that shape decision-making, including a central interaction zone where vehicles must account for potential conflicts and trajectory overlaps, and a broader approach area where speed and heading adjustments begin in anticipation of entering the roundabout. The inner circle represents the main circulation area, while the shaded zones near the exits denote heightened interaction regions where yielding, merging, and exiting are most critical. These regions collectively define the spatial context in which CAVs must coordinate with surrounding agents to ensure safe and efficient navigation, and the key notation is summarized in Table I.
III Safety-Critical Decision Making System
@@ III-A Multi-Agent MDP Formulation for Dual-Lane Roundabout
TABLE I: Summary of Key Notation
Symbol Description
State and Control
$s_{i}=[r_{i},\theta_{i},v_{i},\phi_{i},l_{i}]^{\top}$ State vector of vehicle $i$
$u_{i}=[a_{i},\dot{\phi}_{i},\delta_{i}]^{\top}$ Control input of vehicle $i$
$\mathcal{C},\;\mathcal{H}$ Set of CAVs / HDVs
N,MN,\;M Number of CAVs / HDVs
Safety Assessment
$d(s_{i},s_{j})$ Min. distance between vehicles $i,j$
$d_{\text{safe}}(s_{i},s_{j})$ Adaptive safe distance
$d_{\min}$ Min. base safe distance
$d_{\text{c2b}}(s_{i})$ Min. distance to roundabout boundary
$Q_{\text{risk}}^{\text{cc}},\;Q_{\text{risk}}^{\text{ch}}$ C2C / C2H overall risk metric
$Q_{\text{th}}^{\text{cc}},\;Q_{\text{th}}^{\text{ch}}$ Safety thresholds for C2C / C2H
$C_{ih}$ C2H collision probability
HDV Uncertainty Modeling
$f_{\text{IDM}}(s_{h})$ IDM nominal prediction for HDV $h$
$\boldsymbol{\Sigma}_{h}^{\text{lane}}$ Lane-specific covariance matrix
$\mathbf{M}_{\text{lane}}(l_{h})$ Lane multiplier matrix
$\mathbf{M}_{\text{exit}}(\mathbf{E}_{h})$ Exit proximity multiplier
MCTS and Reward
$N_{n},\;Q_{n}$ Node visit count / cumulative reward
$e_{\text{exp}}$ UCB exploration constant
$K$ Number of MCTS iterations
$\phi_{i}$ Shapley value of agent $i$
$\lambda_{i}^{t}$ Cooperation coefficient
@@
Navigation decisions must also satisfy constraints over $T$:
@@
TABLE II: Safety distance adjustment factors
@@ III-B Hierarchical Safety Assessment for Roundabout Navigation
2) C2C and C2H safety assessment: The adaptive safety distance accounts for multiple risk factors via a base value and adjustment terms, as shown in Table II:
@@
where $d_{\min}$ is the minimum base distance, $\kappa_{v}$ is the velocity scaling factor, $\Delta v_{ij}=v_{i}-v_{j}$ is the relative velocity, $\Delta\phi_{ij}$ is the relative heading angle, $l_{i},l_{j}\in\{0,1\}$ denote lane indices, $v_{\text{ref}}$ is the reference velocity for normalization, $\beta_{k}$ are weighting parameters, and $\mathbb{1}_{\Omega}(\cdot)$ is the indicator function for vehicles in zone $\Omega$. $\mathcal{K}=\{v,\phi,l,z\}$ indexes the adjustment factors in Table II. Fig. 3 visualizes our safety-critical risk assessment framework for CAVs and HDVs navigating roundabouts. Fig. 3(a) presents the foundational distance-based risk function, characterized by an exponential decay curve, which quantifies collision urgency and informs immediate avoidance actions in MCTS decisions. The clearly defined safe distance threshold enables effective and computationally efficient binary safety decisions. Fig. 3(b) provides a spatial risk heat map for a representative roundabout scenario, highlighting high-risk zones at lane-changing and exit-merging points.
@@
3) Lane-Specific HDV Uncertainty Modeling: For C2H interactions, we adopt the IDM [36] as the nominal behavior predictor and embed it within a probabilistic framework to account for behavioral deviations. The lane-specific covariance $\boldsymbol{\Sigma}_{h}^{\text{lane}}$ captures the distinct uncertainty profiles of inner and outer lanes, where drivers exhibit different levels of yielding tendency and speed variability. The uncertainty model is:
@@
The scaling factors are motivated by the geometric asymmetry of roundabout lanes: inner-lane vehicles are constrained by a tighter turning radius that naturally limits lateral and heading deviations, whereas outer-lane vehicles have greater freedom for radial displacement and exit maneuvers [24]. The effectiveness of this lane-specific design over uniform uncertainty is validated in the ablation study (Section V-C).
@@
where $\psi(\hat{s}_{i},\hat{s}_{h})=\mathbb{I}(d(\hat{s}_{i},\hat{s}_{h})<d{\text{safe}}(s_{i},s_{h}))$ is the collision indicator function. The safety level for C2H interactions is:
@@ IV-A Multi-Agent MCTS for Roundabout Navigation
We set the safety thresholds to $Q_{\text{th}}^{\text{cc}}=0.8$ (C2C) and $Q_{\text{th}}^{\text{ch}}=0.6$ (C2H), using a more conservative margin for human-driven interactions due to higher prediction uncertainty. These values are calibrated in preliminary experiments to balance safety and traffic efficiency, and the sensitivity to $d_{\min}$ is further studied in Section V-C. Following standard surrogate safety analysis, we adopt a PET threshold of $1.0$ s [9].
@@
Figure 5: Illustration of the safety-critical MCTS framework. Each node stores the visit count $N_{n}$, value estimate $Q_{n}$, and UCB score. Nodes identified as unsafe (shown in orange) are pruned during the safety validation stage. The green path indicates the backpropagation process.
@@
where $e_{\text{exp}}$ is the exploration constant balancing exploitation and exploration, and $N_{p_{n}}$ is the visit count of the parent node. Fig. 5 illustrates the structure of an MCTS guided by the UCB strategy. Leaf nodes represent terminal or intermediate states, with one marked as an “Unsafe Node” indicating trajectories that violate safety constraints.
@@ IV-D Computational Complexity Analysis
The computational complexity of the proposed multi-agent MCTS-based planning algorithm arises from three main components: joint-action tree expansion, safety validation, and rollout simulation. Tree expansion considers joint actions over $N$ CAVs, each with $|\mathcal{A}|$ possible actions, resulting in a branching factor of $|\mathcal{A}|^{N}$. At each expanded node, safety validation involves pairwise CAV interactions ($O(N^{2})$), CAV-to-HDV checks ($O(NM)$), and lane-level uncertainty evaluations, typically contributing $O(NM\cdot d^{2})$ complexity, where $d$ is the state dimension.
@@
TABLE III: Computational complexity comparison
Method Complexity Operations
Joint exhaustive search $|\mathcal{A}|^{N\cdot H}$ $\approx 10^{85}$
Game-theoretic Nash $\mathcal{O}(|\mathcal{A}|^{NH}\cdot I_{\text{Nash}})$ $10^{87}$
MCTS (no pruning) $K\cdot|\mathcal{A}|\cdot N\cdot H$ $\approx 3.2\times 10^{5}$
MCTS + safety pruning $(1-\rho)\cdot K\cdot|\mathcal{A}|\cdot N\cdot H$ $\approx\mathbf{2.2\times 10^{5}}$
To provide practical runtime context, we report measured per-step computation times from our MATLAB implementation. In our most complex scenario (8-vehicle symmetric intersection), the framework plans for $N=4$ decision-making CAVs while the remaining HDVs are modeled as reactive agents with negligible planning overhead. The average planning time is approximately 58 ms per decision step, well within a typical planning cycle of $\Delta t=0.2$ s. Given the linear scaling with $N$ in our MCTS framework, we estimate the current implementation can support approximately N=6N=688 planning agents in real time. As shown in Table III, safety pruning reduces the effective search space by approximately $1.4\times$ compared to unpruned MCTS, while UCB-guided selection further concentrates search effort on promising branches. A C++ implementation with parallelized rollouts would further extend this scalability envelope, left for future work.
(a) Illustration of MCTS simulation at a dual-lane roundabout.
@@ V EXPERIMENTAL EVALUATION
(a) Illustration of MCTS simulation at a dual-lane roundabout.
@@ V-A Case 1: Dual-lane Roundabout (ROP = 100%)
Methods Average Speed Average Trajectory Collision
(m/s) Deviation (m) Rate (%)
Stackelberg $5.22\pm 1.34$ $0.89\pm 0.84$ $16.0$
Nash $5.16\pm 1.01$ $0.74\pm 0.59$ $11.0$
Baseline $5.44\pm 1.10$ $0.65\pm 0.47$ $13.0$
$d_{\max}=3$ $5.27\pm 1.03$ $0.57\pm 0.38$ $4.0$
$d_{\max}=8$ $\mathbf{5.31\pm 0.89}$ $\mathbf{0.45\pm 0.39}$ $\mathbf{0.0}$
@@
Fig. 7(b) presents a quantitative assessment of safety performance based on Post-Encroachment Time (PET) distributions. The proposed method with a maximum tree depth of $d_{\text{max}}=8$ delivers the best outcomes, with no instances of PET falling below the critical threshold of $1.0$ second. In comparison, the baseline approach exhibits $8.8\%$ violations, while the Stackelberg and Nash methods perform worse, with violation rates of $26.1\%$ and $13.2\%$, respectively. These results underscore the effectiveness of our framework, which leverages dynamic risk evaluation and adaptive safety constraints to proactively manage both current and anticipated interactions.
Table IV demonstrates the superior performance of our proposed method with $d_{\max}=8$, which achieves the lowest collision rate (0.0%) and the smallest trajectory deviation ($0.45\pm 0.39$,m), while maintaining competitive driving speed ($5.31\pm 0.89$,m/s). In contrast, Stackelberg and Nash baselines exhibit substantially higher collision rates (16.0% and 11.0%), indicating their limited robustness in complex interaction scenarios. Increasing $d_{\max}$ consistently improves safety, confirming the benefit of incorporating larger safety margins in roundabout coordination.
@@
Methods Average Speed Average Trajectory Collision
(m/s) Deviation (m) Rate (%)
Stackelberg $5.25\pm 1.31$ $1.14\pm 1.06$ $33.0$
Nash $5.12\pm 1.24$ $0.85\pm 0.68$ $29$
Baseline $4.83\pm 0.82$ $0.81\pm 0.65$ $17$
$d_{\max}=3$ $5.16\pm 0.95$ $0.76\pm 0.62$ $5.0$
$d_{\max}=8$ $\mathbf{5.23\pm 0.94}$ $\mathbf{0.63\pm 0.59}$ $\mathbf{2.0}$